OpenStudy (anonymous):

find the value of n such that x^2-11x+n is a perfect square trinomial

OpenStudy (anonymous):

$$(x-c)^2=x^2-2cx+c^2$$Here we have \(2c=11\) so \(c=11/2\) and \(n=c^2=121/4\)

OpenStudy (anonymous):

\[\sqrt{k-4=11}\]